OpenStudy (anonymous):

Thelma wants to find the equation of a line that passes through the point (2, 2) and is perpendicular to the line y=1/2x+5. What is the equation?

OpenStudy (btaylor):

use this equation: \[y-y_{1}=m(x-x_{1})\] Plug in your known values. m=-2, since the perpendicular slope is the negative reciprocal of the original slope. You get: \[y-2 = -2(x-2)\] Solve for y: \[y = -2x+6\]

OpenStudy (anonymous):

thanks....i will have this yet!